A metal sphere, when suspended in a constant temperature enclosure, cools from 80 0C to 70 0C in 5 minutes and cool from 70 0C to 62 0C in the next five minutes. Calculate the temperature of the enclosure.  


1
Expert's answer
2020-12-31T08:07:58-0500

Apply Newton's law of cooling, where "\theta_0" is the temperature of surroundings:


"\frac{\Delta\theta}{\Delta t}=k(\theta-\theta_0)."


For the first case, we have


"\frac{\theta_1-\theta_2}{\Delta t}=k\bigg(\frac{\theta_1+\theta_2}{2}-\theta_0\bigg),\\\space\\
\frac{80-70}{5}=k\bigg(\frac{80+70}{2}-\theta_0\bigg)."

For the second case, by analogy


"\frac{70-62}{5}=k\bigg(\frac{70+62}{2}-\theta_0\bigg)."

Divide one by another:


"\frac{80-70}{70-62}=\frac{75-\theta_0}{66-\theta_0},\\\space\\
\theta=30\text{°C}."